OpenStudy (anonymous):

find all the other functions of θ if it is in the quadrant indicated . 1. SIN θ =4/5 (1ST QUADRANT) 2. COS θ = -12/13 (3RD QUADRANT) 3. COT θ = 4/3 (3RD QUAD) 4. SEC θ = -5/3 (3RD QUAD) 5. SIN θ = 7/25 (2ND QUAD)

OpenStudy (anonymous):

You mostly have to use \[\sin^2(x) + \cos^2(x) = 1\] to find the missing one, as well as \[\tan(x) = \sin(x)/\cos(x)\]

OpenStudy (anonymous):

so how do i know what's the answer i be looking on google but cant find any idea

OpenStudy (anonymous):

I gave you the process with which you can find the missing one (simply replace sinx = 4/5 and solve for cos(x), for instance). You obviously have to know the definitions of each trig function too...

OpenStudy (king):

cosx=3/5

OpenStudy (king):

tanx=4/3

OpenStudy (anonymous):

I will give you one example, but you must work out the rest yourself. sinx = 4/5, and we are in the first quadrant. Using \[\sin^2(x) + \cos^2(x) = 1 <=>(4/5)^2 + \cos^2(x) = 1 <=> \cos(x) = +/- \sqrt{1 -(4/5)^2}\] = +/- 3/5 Because we are in the first quadrant, the cosine is positive, so the solution is + 3/5

OpenStudy (king):

cos

OpenStudy (king):

yeah + or - 3/5

OpenStudy (anonymous):

if it is in 1st quadrant it is positive right?

OpenStudy (king):

yeah but just in case if quad not given then u should write +or-

OpenStudy (king):

ok so hav u got it?

OpenStudy (anonymous):

in the end i will use phytagorean theorem? then can i use calculator?

OpenStudy (king):

no need to use calc. please dont use it is very bad to use a calculator

OpenStudy (anonymous):

The calculator is not needed for these, only the definitions of each trigonometric function, and the 2 identities which I gave you

OpenStudy (anonymous):

Please read my posts carefully, it feels like you're skipping them.

OpenStudy (king):

yeah ringlum is right and his sols. are also right please read them

OpenStudy (king):

and btw if there are any Arsenal F.C fans please join my group 'Arsenal'

OpenStudy (anonymous):

so if i feel no.1 and no.2 solution is the same?

OpenStudy (anonymous):

because sin^2 (x) + cos ^2(x) = 1 right?

OpenStudy (king):

he just used the identities of his first post to write the sol. in his second post

OpenStudy (king):

yeah

OpenStudy (king):

waitin sis

OpenStudy (anonymous):

(what I wrote applies to the first, number 1, you must do the same process or similar for each of the other ones)

OpenStudy (anonymous):

-12/13 + sin^ (x) = 1 sin x =